OpenStudy (anonymous):

What is the slope for the line tangent to the circle x2 + y2 = 8 at the point (2, -2)?

OpenStudy (dumbcow):

slope = dy/dx at point (2,-2) differentiate circle equation and solve for dy/dx \[\rightarrow 2x +2y \frac{dy}{dx} = 0\] \[\rightarrow \frac{dy}{dx} = -\frac{x}{y}\] plug in point \[\frac{dy}{dx} = \frac{-2}{-2} = 1\]